Grade 11Mechanics

A constant force of 40 N is applied tangentially to the rim of a wheel of radlus 20 cm. The wheel has moment of inertia30 kg m2. Ten the work done on wheel in 4 second is

Nitin Kumar
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
From the centre of mass,
F * R = I*a/R
40 * 20/100 = 30*a*100/20
a = 8*2/300 => 16/300 => 8/150 => 4/75 m/s^2
v = a*t => 4*4/75 => 16/75 m/s
Power = 40 * 16/75 Watt
Work = Power x time > 40*16*4/75 Joule =>  34.13 Joules